5.3 gms of NA2CO3  ARE DISSOLVED  IN 100 ML OF WATER,THEN NORMALITY OF THE SOLUTION IS


 

Grade:

3 Answers

Sunil Kumar FP
10 years ago
normality=wt/eqwt×v(l)
here eq wt=mw/2=106/2=53
therefore ,normality=5.3/53×.1=1N

thanks and regard
sunil kumar
ask iitian
KARTHIK
10 years ago
Normality=wt/eqwt×1000/v(ml) (5.3÷106)(1000÷100=0.5
KARTHIK
10 years ago
Normality=wt/eqwt×1000/v(ml) (5.3÷106)(1000÷100)=0.5